OpenStudy (skullcandy229):

Find the sum. In your final answer, include your calculations. -3.45 + 5.1

OpenStudy (ninjaslice):

1.65

OpenStudy (ninjaslice):

-3.45 + 5.1 = 1.65 DUE TO 5.1 - 3.45 = 1.65

OpenStudy (trojanpoem):

Take it as arts: -3 - 0.45 + 5 + 0.1 -3 + 5 = 2 -0.45 + 0.10 = -0.35 final answer is 2 - 0.35 = 1. (100 - 35) 1.65

OpenStudy (skullcandy229):